OpenStudy (anonymous):

which of the following represents the slope of the equation 57x-95y=0 5/3 or 3/2 or 2/3 or 3/5

OpenStudy (anonymous):

\[y= \frac{57x}{95}\]

OpenStudy (anonymous):

in the form y=mx +b so gradient = m = 57/95 = 3/5

OpenStudy (anonymous):

thanks